How much pure alcohol should be added to 400 ml. of 15% strength to make its strength 32%?

Answer by mathmandotcom(31)

.15(400) +1x=.32(400+x): 1=100% x=amount of prue alcohol needed
60+x=120+.32x: subtract 60 from both sides
x=60+.32x: subtract .32x from both sides
.68x=60; divide by .68
x=60/.68; simplify
x=88.235ml at least